OpenStudy (anonymous):

A box-and-whisker plot is created for the following data set: 2, 4, 6, 8, 10, 12, 14. Which of the data items below could be increased by 1 without changing the plot?

OpenStudy (anonymous):

I'm guessind its 2?

OpenStudy (amistre64):

2 changes the range

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

first and last; and q1,q2,q3 have to remain unchanged

OpenStudy (amistre64):

so there are 5 elements that cant change

OpenStudy (anonymous):

the choices are 2,4,6,8

OpenStudy (amistre64):

well, 8 is the median so that has to remain 2 and 14 are the ends so they have to remain the same define the q1 and q3

OpenStudy (anonymous):

4 and 6?

OpenStudy (amistre64):

there are 7 elements q1 = 7/4 = 2.xx; or 3rd number q3 = 3*7/4 = 5.xx; or the 6th number

OpenStudy (amistre64):

2, 4, 6, 8, 10, 12, 14. ^ ^ ^ ^ ^ those have to remain the same

OpenStudy (anonymous):

so the answer is 4

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

YAY thanks!

OpenStudy (amistre64):

let me chk my senility first

OpenStudy (amistre64):

7/4 is not 2.xx its 1.xx so the 2nd potiion is q1

OpenStudy (amistre64):

2, 4, 6, 8, 10, 12, 14. ^ ^ ^ ^ ^ thats better

OpenStudy (amistre64):

id go with 6 :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

next time i try to divde 7 by 4 and get over 2 .... let me know ;)

OpenStudy (anonymous):

ok i will